The world’s smallest mammal is the bumblebee bat. Such bats are roughly the size of a large bumblebee. A random sample of 130 of these bats had an average weight of 1.7 grams with a standard deviation of 0.22 grams. Construct a 95% confidence interval estimate of their mean weight
Answer #1

Here we have: x = 1.7 S= 0.22 n = 130 a = 1-0.95 = 0.05 (95% confidence level) df = n-1 = 130-1 = 129 Critical value: tarit=

= 1.7+ 0.038 = 1.662 <μ<1.738